During the first 20 mi of a 50 mi bicycle race. Roger's average speed was 16 mi/h. what must his average speed be during the remainder of the race if he is to finish the race in less than 2.5h?
Answer by mananth(12270)

During the first 20 mi of a 50 mi bicycle race. Roger's average speed was 16 mi/h. what must his average speed be during the remainder of the race if he is to finish the race in less than 2.5h?
First 20 mi. 16 mph
time = 20/16 = 5/4 hours.
..
he has to finish in less than 2.5 hours = 5/2 hours
Balance time = 5/2 - 5/4 = 10-5 /4= 5/4 hours
.
He has to ride 30 more miles.
in less than 5/4 hours
speed > distance / time
speed > 30/(5/4)
speed > 30*4/5
speed > 24 mph
he has to ride at speed > 24 mph